### v1

# Multi-virtual braid groups

Abstract.
L. Kauffman (2024) introduced multi-virtual and symmetric multi-virtual braid groups, which are generalizations of the virtual braid group. We introduce multi-virtual pure and multi-virtual semi-pure braid groups, which are normal subgroups of index $n!$. We give a set of generators and defining relations for these groups, show that multi-virtual (symmetric multi-virtual) braid group is a semi-direct products of multi-virtual pure (symmetric multi-virtual pure) braid group and symmetric group. Also, we introduce multi-welded and multi-unrestricted braid groups and examines structure of three-strand 2-virtual braid group and some its subgroups and quotients. The paper concludes by outlining open problems and suggesting avenues for future research in this area.
Key words and phrases:
Braid group, pure braid group, Multi-virtual braid group, symmetric multi-virtual braid group.
2010 Mathematics Subject Classification:
20E07, 20F36, 57K12
1. Introduction
The braid groups, $B_{n}$, $n\geq 1$ form a basis of classical knot theory. During some last decades were introduced and intensively studied different generalizations of $B_{n}$. For example, virtual braid group $VB_{n}$, welded braid group $WB_{n}$, singular braid group $SB_{n}$ and some other (see [3, 10, 15] and references therein). These groups form basis of the virtual knot theory, the welded knot theory and the singular knot theory respectively. Also, it is possible to construct some quotient of $VB_{n}$. On this way, were defined flat virtual braid group $FVB_{n}$, unrestricted virtual braid group $UVB_{n}$ [5] and some others.
All these groups $VB_{n}$, $WB_{n}$, $SB_{n}$, $FVB_{n}$, and $UVB_{n}$ have two types of generators that corresponds to two types of crossings in the corresponding knot theories. It is natural to call these groups by 2-sort braid-like groups. There are braid-like groups, which have three types of generators. For example, singular virtual braid groups [6, 7], twisted virtual braid group [2]. We shall call them by 3-sort braid-like groups. Also, singular twisted virtual braid groups are the 4-sort braid-like groups studied in [20].
L. Kauffman [14] introduced multi-virtual braid groups. These groups are a generalization of the virtual braid group and contains many types of virtual generators. Multi-virtual braid groups correspond to multi-virtual knot theory with many types of virtual crossings. Defining relations in the multi-virtual braid groups are defined by detour moves which allowed with respect to some types of crossings, but are forbidden between other crossings.
The braid group $B_{n}$ has an epimorphism onto the symmetric group $S_{n}$, under which the generator $\sigma_{i}$ goes to the transposition $(i,i+1)$. The kernel of this epimorphism is said to be a pure braid group. By analogy, the virtual pure braid group $VP_{n}$ is the kernel of the epimorphism of the virtual braid group $VB_{n}$ onto $S_{n}$ that maps, for each $i$, $\sigma_{i}$ and $\rho_{i}$ to the $\rho_{i}$. Here $\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle\cong S_{n}$.
In this paper we are studying the following questions. Find generators and defining relations for pure and semi-pure multi-virtual braid groups. Define some quotients of these groups and find their generators and defining relations.
The paper is organized as follows. In Section 2, we recall some known definitions and known facts from braid theory. In Subsections 2.1 and 2.2 one can finds definitions and some known facts from classical and virtual braid theories. In particular, we give definitions and presentations for virtual pure and virtual semi-pure braid groups. In Subsection 2.3, we recall multi-virtual knot theory in brief.
In Section 3, we recall the Kauffman’s definitions of multi-virtual braid group and symmetric multi-virtual braid group. Further, we introduce multi-virtual pure braid group and multi-virtual semi-pure braid group. We give sets of generators and defining relations for these groups. Similar results we give for the symmetric multi-virtual braid group. In Section 4, we introduce multi-welded braid group and multi-unrestricted braid groups, we find their set of generators and defining relation. Also, we defined pure and semi-pure subgroups in these groups. In Section 5, we find presentations for 2-virtual braid group with 3-strand subgroups and presentations for some subgroups and quotients.
In the end of the paper, we formulate some open problems and suggest directions for further research.
2. Basic definitions
In this section we recall some known definitions which can be found in [1, 11, 19].
2.1. Braid group
The braid group $B_{n}$, $n\geq 2$, on $n$ strands can be defined as a group generated by $\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1}$ with the defining relations
(1) $\sigma_{i}\,\sigma_{i+1}\,\sigma_{i}=\sigma_{i+1}\,\sigma_{i}\,\sigma_{i+1},~{}~{}~{}i=1,2,\ldots,n-2,$
(2) $\sigma_{i}\,\sigma_{j}=\sigma_{j}\,\sigma_{i},~{}~{}~{}|i-j|\geq 2.$
The geometric interpretation of $\sigma_{i}$ and its inverse $\sigma_{i}^{-1}$ are depicted in the Figure 1.
Figure 1. The geometric interpretation of $\sigma_{i}$ and $\sigma_{i}^{-1}$
2.2. Virtual braid group
The virtual braid group $VB_{n}$ was introduced in [13]. This group is generated by the braid group $B_{n}$ and the symmetric group $S_{n}=\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle$ with the following relations:
$\sigma_{i}\sigma_{i+1}\sigma_{i}$ $=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ $i=1,2,\ldots,{n-2},$
$\sigma_{i}\sigma_{j}$ $=\sigma_{j}\sigma_{i}$ $|i-j|\geq 2,$
$\rho_{i}^{2}$ $=e$ $i=1,2,\ldots,{n-1},$
$\rho_{i}\rho_{j}$ $=\rho_{j}\rho_{i}$ $|i-j|\geq 2,$
$\rho_{i}\rho_{i+1}\rho_{i}$ $=\rho_{i+1}\rho_{i}\rho_{i+1}$ $i=1,2,\ldots,{n-2},$
$\sigma_{i}\rho_{j}$ $=\rho_{j}\sigma_{i}$ $|i-j|\geq 2,$
$\rho_{i}\rho_{i+1}\sigma_{i}$ $=\sigma_{i+1}\rho_{i}\rho_{i+1}$ $i=1,2,\ldots,{n-2}.$
[…]
(3) $\lambda_{ij}\lambda_{kl}=\lambda_{kl}\lambda_{ij},$
(4) $\lambda_{ik}\lambda_{jk}\lambda_{ij}=\lambda_{ij}\lambda_{jk}\lambda_{ik},$
where distinct letters stand for distinct indices. The generators of $VP_{n}$ can be expressed in terms of the generators of $VB_{n}$ by the formulas
$\lambda_{i,i+1}$ $=\rho_{i}\,\sigma_{i}^{-1},$
$\lambda_{i+1,i}$ $=\rho_{i}\,\lambda_{i,i+1}\,\rho_{i}=\sigma_{i}^{-1}\,\rho_{i}$
for $i=1,2,\ldots,n-1$, and
$\lambda_{i,j}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,\lambda_{i,i+1}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
$\lambda_{j,i}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,\lambda_{i+1,i}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1}$
for $11\leq i<j-1\leq n-$ 1.
We have decomposition $VB_{n}=VP_{n}\rtimes S_{n}$ and $S_{n}$ acts on $VP_{n}$ by the rules
Lemma 2.1 ([3]).
Let $a$ be an element of $\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle$ and $\bar{a}$ is its image in $S_{n}$ under the isomorphism $\rho_{i}\mapsto(i,i+1)$, $i=1,2,\ldots,n-1$, then for any generator $\lambda_{ij}$ of $VP_{n}$ the following holds
$a^{-1}\lambda_{ij}a=\lambda_{(i)\bar{a},(j)\bar{a}},$
where $(k)\bar{a}$ is the image of $k$ under the action of the permutation $\bar{a}$.
[…]
Example 2.2.
$VP_{3}$ is generated by elements
$\lambda_{12},~{}~{}\lambda_{21},~{}~{}\lambda_{13},~{}~{}\lambda_{23},~{}~{}\lambda_{31},~{}~{}\lambda_{32},$
and is defined by six relations
$\lambda_{ki}\lambda_{kj}\lambda_{ij}=\lambda_{ij}\lambda_{kj}\lambda_{ki},~{}~{}i,j,k\in\{1,2,3\}.$
It contains
$FVP_{3}=\langle\lambda_{12},\lambda_{13},\lambda_{23}~{}|~{}\lambda_{12}\lambda_{13}\lambda_{23}=\lambda_{23}\lambda_{13}\lambda_{12}\rangle$
that is a group with one defining relation. This group is the flat virtual pure braid group and is the quotient of $VP_{3}$ by relations $\lambda_{ij}=\lambda_{ji}^{-1}$.
It is possible to define another epimorphism $\psi\colon VB_{n}\to S_{n}$ as follows:
$\psi(\sigma_{i})=e,\;\psi(\rho_{i})=\rho_{i},\;i=1,2,\dots,n-1.$
Let $VH_{n}$ denote the normal closure of $B_{n}$ in $VB_{n}$. It is evident that $\ker\psi$ coincides with $VH_{n}$. Let us define elements:
$x_{i,i+1}=\sigma_{i},~{}~{}x_{i+1,i}=\rho_{i}\sigma_{i}\rho_{i}=\rho_{i}x_{i,i+1}\rho_{i},$
for $i=1,2,\ldots,n-1$, and
$x_{i,j}=\rho_{j-1}\cdots\rho_{i+1}\sigma_{i}\rho_{i+1}\cdots\rho_{j-1},$
$x_{j,i}=\rho_{j-1}\cdots\rho_{i+1}\rho_{i}\sigma_{i}\rho_{i}\rho_{i+1}\cdots\rho_{j-1},$
for $11\leq i<j-1\leq n-$ 1.
The group $VH_{n}$ admits a presentation with the generators $x_{k,\,l},$ $1\leq k\neq l\leq n$, and the defining relations:
(5) $x_{i,j}\,x_{k,\,l}=x_{k,\,l}\,x_{i,j},$
(6) $x_{i,k}\,x_{k,j}\,x_{i,k}=x_{k,j}\,x_{i,k}\,x_{k,j},$
[…]
Lemma 2.3 ([4]).
Let $a$ be an element of $\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle$ and $\bar{a}$ is its image in the symmetric group $S_{n}$ under the isomorphism $\rho_{i}\mapsto(i,i+1)$, $i=1,2,\ldots,n-1$, then for any generator $x_{ij}$ of $VH_{n}$ the following holds
$a^{-1}x_{ij}a=x_{(i)\bar{a},(j)\bar{a}},$
where $(k)\bar{a}$ is the image of $k$ under the action of the permutation $\bar{a}$.
2.3. Knot theory and multi-virtual knot theory
The famous Reidemeister’s theorem [22] says that the set of local moves $(r1)-(r3)$ of link’s diagrams, shown in the first row in Figure 2 generate isotopy of knots and links. Reidemeister showed that if we have two knots (links) in 3-dimensional space, then they are ambient isotopic if and only if corresponding diagrams for them can be obtained, one from the other, by a finite sequence of moves $(r1)-(r3)$.
Figure 2. Reidemeister and virtual Reidemeister moves
A virtual link represents a natural combinatorial generalization of a classical link (see [13]). A virtual diagram (or a diagram of a virtual link) is the image of an immersion of a framed 4-valent graph in $\mathbb{R}^{2}$ with a finite number of intersections of edges. Moreover, each intersection is a transverse double point which we call a virtual crossing and mark by a small circle, and each vertex of the graph is endowed with the classical crossing structure (with a choice for underpass and overpass specified). The vertices of the graph are called classical crossings or just crossings.
Two virtual link diagrams are virtually equivalent if one diagram can be transformed into the other by a sequence of classical and virtual Reidemeister moves as shown in Figure  2. The virtual Reidemeister moves are equivalent to a single move, the detour move, which is executed by selecting a segment of a component of the link diagram that contains no classical crossings (see Figure 3). After removing this segment, we may insert a new segment with no triple points and any double points result in a new virtual crossing. A virtual link is an equivalence class of virtual diagrams modulo the Reidemeister moves referring to classical crossings and the detour moves.
Figure 3. Detour move for a virtual crossings of type 00
L. Kauffman [14] introduced multi-virtual knot theory with many types of virtual crossings. A diagram of multi-virtual link is analogous to a diagram of a virtual link, but in this case we have many types of virtual crossings. A multi-virtual link is an equivalence class of multi-virtual diagrams modulo the Reidemeister moves referring to classical crossings and the detour moves, in which we transform an arc, which contains virtual crossings only one type (see Figure 3, in which we denote this type of virtual crossings by 00). Remark that the detour moves with over types of virtual crossings are forbidden (see Figure 4, where $\alpha\not=0$).
Figure 4. Forbidden detour move for a virtual crossings of type $\alpha\not=0$
3. Multi-virtual braid group
In this section, we recall (see [14]) definitions of the multi-virtual and the symmetric multi-virtual braid groups. The motivation for introduction of these groups is the multi-virtual knot theory (see Subsection 2.3), which is a generalization of the classical knot theory.
Let $k\geq 1$, $k$-multiple virtual braid group on $n$ strands, $M_{k}VB_{n}$ is a group which is generated by elements
$\sigma_{i},\rho^{(\alpha)}_{i},~{}~{}i=1,2,\ldots,n-1,~{}\alpha=0,1,\ldots,k-1,$
and is defined by the following types of relations:
I. Involutivity of generators
$\left(\rho^{(\alpha)}_{i}\right)^{2}=e,~{}~{}i=1,2,\ldots,n-1,~{}\alpha=0,1,\ldots,k-1.$
II. For commutativity
– homogeneous
$\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i},\,\,\rho_{i}^{(\alpha)}\rho_{j}^{(\alpha)}=\rho_{j}^{(\alpha)}\rho_{i}^{(\alpha)},~{}~{}|i-j|\geq 2;$
– mixed ($\alpha\not=\beta$)
$\sigma_{i}\rho_{j}^{(\alpha)}=\rho_{j}^{(\alpha)}\sigma_{i},\,\,\rho_{i}^{(\alpha)}\rho_{j}^{(\beta)}=\rho_{j}^{(\beta)}\rho_{i}^{(\alpha)},~{}~{}|i-j|\geq 2.$
III. Braid relations
– homogeneous
$\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1},\,\,\,\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)},~{}~{}i=1,2,\ldots,{n-2};$
– mixed ($\beta=1,2,\ldots,k-1$)
$\sigma_{i}\rho_{i+1}^{(0)}\,\rho_{i}^{(0)}=\rho_{i+1}^{(0)}\,\rho_{i}^{(0)}\sigma_{i+1},$
$\rho_{i}^{(0)}\,\rho_{i+1}^{(0)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(0)}\,\rho_{i+1}^{(0)},~{}~{}~{}~{}i=1,2,\ldots,{n-2}.$
Remark 3.1.
Relations of the following forms are forbidden
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $10\leq\alpha\leq k-$ 1;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $10\leq\alpha\leq k-$ 1;
$F3\colon\rho_{i}^{(0)}\,\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}\,\rho_{i+1}^{(0)}$,   $\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}\,\rho_{i+1}^{(\gamma)}$,
$\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\gamma)}\,\rho_{i+1}^{(\gamma)}$,   $10<\gamma<\beta\leq k-$ 1.
Remark 3.2.
From the mixed braid relations it follows the relations
$\sigma_{i+1}\rho_{i}^{(0)}\,\rho_{i+1}^{(0)}=\rho_{i}^{(0)}\,\rho_{i+1}^{(0)}\sigma_{i}$
and the relations
$\rho_{i+1}^{(0)}\,\rho_{i}^{(0)}\,\rho_{i+1}^{(\beta)}=\rho_{i}^{(\beta)}\,\rho_{i+1}^{(0)}\,\rho_{i}^{(0)},~{}~{}~{}~{}i=1,2,\ldots,{n-1}.$
If $k=1$, then $M_{k}VB_{n}$ is the virtual braid group which is denoted by $VB_{n}$.
Proposition 3.3.
  1. (1)
    $VB_{n}\leq M_{k}VB_{n}$.
  2. (2)
    If $k\geq 2$, then $FVB_{n}\leq M_{k}VB_{n}$.
  3. (3)
    For $k\geq 1$, $M_{k}VB_{n}\leq M_{k+1}VB_{n}$.
Proof.
  • (1)
    Define a map $\psi_{1}\colon M_{k}VB_{n}\to VB_{n}$ such that $\sigma_{i}\mapsto\sigma_{i}$, $\rho_{i}^{(0)}\mapsto\rho_{i}$ and $\rho_{i}^{(\alpha)}\mapsto e$. This map is a homomorphism. Define a section of $\psi_{1}$, an inclusion map $\iota\colon VB_{n}\hookrightarrow M_{k}VB_{n}$ such that $\psi_{1}\circ\iota$ is the identity map on $VB_{n}$. The map $\iota\colon VB_{n}\to M_{k}VB_{n}$ is a well defined homomorphism. Since the homomorphism $\psi_{1}$ has a section $\iota$ then the image $VB_{n}$ is a subgroup of $M_{k}VB_{n}$.
  • (2)
    Define a map $\psi_{2}\colon M_{k}VB_{n}\to FVB_{n}$ such that $\sigma_{i}\mapsto e$, $\rho_{i}^{(0)}\mapsto\rho_{i}$, $\rho_{i}^{(1)}\mapsto c_{i}$ and $\rho_{i}^{(\alpha)}\mapsto e$ for $\alpha\neq 0,1$. This map is a homomorphism. Define a section of $\psi_{2}$, an inclusion map $\iota\colon FVB_{n}\hookrightarrow M_{k}VB_{n}$ $c_{i}\mapsto\rho^{(1)}_{i}$ and $\rho_{i}\mapsto\rho_{i}^{(0)}$ such that $\psi_{2}\circ\iota$ is the identity map on $FVB_{n}$. The map $\iota\colon FVB_{n}\to M_{k}VB_{n}$ is a well defined homomorphism. Since the homomorphism $\psi_{2}$ has a section $\iota$ then the image $FVB_{n}$ is a subgroup of $M_{k}VB_{n}$.
  • (3)
    Define a map $\psi_{3}\colon M_{k+1}VB_{n}\to M_{k}VB_{n}$ such that $\sigma_{i}\mapsto\sigma_{i}$, $\rho_{i}^{(\alpha)}\mapsto\rho^{(\alpha)}_{i}$ for $\alpha\neq k$, and $\rho_{i}^{(k)}\mapsto e$. This map is a homomorphism. Define a section of $\psi_{3}$, an inclusion map $\iota:M_{k}VB_{n}\hookrightarrow M_{k+1}VB_{n}$ such that $\psi_{3}\circ\iota$ is the identity map on $M_{k}VB_{n}$. The map $\iota\colon M_{k}VB_{n}\to M_{k+1}VB_{n}$ is a well defined homomorphism. Since the homomorphism $\psi_{3}$ has a section $\iota$ then the image $M_{k}VB_{n}$ is a subgroup of $M_{k+1}VB_{n}$.
3.1. Multi-virtual pure braid group
Further we shall assume that $\rho_{i}=\rho_{i}^{(0)}$. Since, $S_{n}=\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle\leq VB_{n}$ and $VB_{n}\leq M_{k}VB_{n}$, hence, $S_{n}$ is a subgroup of $M_{k}VB_{n}$.
Let us define a map
$\varphi_{n,k}\colon M_{k}VB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\varphi_{n,k}$ is the multi virtual pure braid group $M_{k}VP_{n}$.
To find generators and defining relations for $M_{k}VP_{n}$, we use the Reidemeister-Schreier method (see, for example [18, Chapter 2.3]). As a Schreier set of coset representation of $M_{k}VP_{n}$ in $M_{k}VB_{n}$ we take the same set $\Lambda_{n}$, which is used in calculation of generators and relations for the group $VP_{n}$,
$\Lambda_{n}=\left\{\prod\limits_{k=2}^{n}m_{k,j_{k}}~{}|~{}1\leq j_{k}\leq k\right\}$
where $m_{kl}=\rho_{k-1}\rho_{k-2}\cdots\rho_{l}$ for $l<k$ and $m_{kl}=1$ in the other cases.
Let us define elements that can be expressed in terms of the generators of $M_{k}VB_{n}$ by the formulas
$\lambda_{i,i+1}^{(0)}$ $=\rho_{i}\,\sigma_{i}^{-1},$
$\lambda_{i+1,i}^{(0)}$ $=\rho_{i}\,\lambda_{i,i+1}^{(0)}\,\rho_{i}=\sigma_{i}^{-1}\,\rho_{i},$
$\lambda_{i,i+1}^{(\beta)}$ $=\rho_{i}\,\rho_{i}^{(\beta)},~{}~{}1\leq\beta\leq k-1,$
for $i=1,2,\ldots,n-1$, and
$\lambda_{i,j}^{(0)}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,\lambda_{i,i+1}^{(0)}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
$\lambda_{j,i}^{(0)}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,\lambda_{i+1,i}^{(0)}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
$\lambda_{i,j}^{(\beta)}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,\lambda_{i,i+1}^{(\beta)}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
for $11\leq i<j-1\leq n-1,~{}~{}1\leq\beta\leq k-$ 1.
Lemma 3.4.
Let $a$ be an element of $\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle$ and $\bar{a}$ is its image in $S_{n}$ under the isomorphism $\rho_{i}\mapsto(i,i+1)$, $i=1,2,\ldots,n-1$, then for any generator $\lambda_{ij}^{(\alpha)}$ of $M_{k}VP_{n}$ the following holds
$a^{-1}\lambda_{ij}^{(\alpha)}a=\lambda_{(i)\bar{a},(j)\bar{a}}^{(\alpha)},$
where $(k)\bar{a}$ is the image of $k$ under the action of the permutation $\bar{a}$.
Theorem 3.5.
$M_{k}VP_{n}$ admits a presentation with the generators
$\lambda^{(0)}_{ij},~{}~{}1\leq i\neq j\leq n,$
$\lambda^{(\beta)}_{ij},~{}~{}1\leq i<j\leq n,~{}~{}1\leq\beta\leq k-1,$
and the following relations:
(7) $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\gamma)}=\lambda_{kl}^{(\gamma)}\lambda_{ij}^{(\alpha)},~{}~{}0\leq\alpha\leq\gamma\leq k-1,$
(8) $\lambda_{ik}^{(0)}\lambda_{jk}^{(0)}\lambda_{ij}^{(0)}=\lambda_{ij}^{(0)}\lambda_{jk}^{(0)}\lambda_{ik}^{(0)},~{}~{}1\leq i,j,k\leq n,$
(9) $\lambda_{ik}^{(\beta)}\lambda_{jk}^{(\beta)}\lambda_{ij}^{(\beta)}=\lambda_{ij}^{(\beta)}\lambda_{jk}^{(\beta)}\lambda_{ik}^{(\beta)},~{}~{}1\leq i<j<k\leq n,~{}~{}1\leq\beta\leq k-1,$
where distinct letters stand for distinct indices.
Proof.
Define the map ¯$\bar{}\colon M_{k}VB_{n}\to\Lambda_{n}$ which takes an element $w\in M_{k}VB_{n}$ to its representative $\overline{w}$ from $\Lambda_{n}$. In this case the element $w\overline{w}^{-1}$ belongs to $M_{k}VP_{n}$. By Theorem 2.7 of [18] the group $M_{k}VP_{n}$ is generated by
$s_{\lambda,a}=\lambda a\cdot(\overline{\lambda a})^{-1},~{}~{}~{}\lambda\in\Lambda_{n},~{}~{}a\in\{\rho^{(\alpha)}_{1},\ldots,\rho^{(\alpha)}_{n-1},\sigma_{1},\ldots,\sigma_{n-1}\},$
where $\alpha=0,1,\ldots,k-1$.
If $\alpha=0$, then all $s_{\lambda,\rho^{(0)}_{i}}=e$ and $s_{\lambda,\sigma_{i}}=\lambda(s_{e,\sigma_{i}})\lambda^{-1}=\lambda(\sigma_{i}\rho_{i})\lambda^{-1}=\lambda(\lambda_{i,i+1}^{-1})\lambda^{-1}$, which is equal to some $\lambda_{kl}$, by Lemma 2.1. These calculations are done in Theorem 1 [3].
Now, consider the generators such that $\alpha\neq 0$.
$s_{\lambda,\rho^{(\alpha)}_{i}}=\lambda(s_{e,\rho^{(\alpha)}_{i}})\lambda^{-1}.$
Since $s_{e,\rho^{(\alpha)}_{i}}=\rho^{(\alpha)}_{i}\rho^{(0)}_{i}$, $s_{\lambda,\rho^{(\alpha)}_{i}}=\lambda(\rho^{(\alpha)}_{i})\lambda^{-1}$, which is equal to some $\lambda^{(\alpha)}_{kl}$ by Lemma 3.4. Therefore, generators of the group $M_{k}VP_{n}$ are
$\lambda^{(0)}_{kl},\text{ for }1\leq k\neq l\leq n\text{ and }\lambda^{(\beta)}_{ij},~{}~{}1\leq i<j\leq n,~{}~{}1\leq\beta\leq k-1.$
To find the defining relations of $M_{k}VP_{n}$, we define a rewriting process $\tau$. It helps to rewrite a word $u$ to $\tau(u)$, where $u$ is written in the generators of $M_{k}VB_{n}$ but represents an element of $M_{k}VP_{n}$ and $\tau(u)$ is a word written in the generators of $M_{k}VP_{n}$. Let us associate to reduce word
$u=a_{1}^{\epsilon_{1}}a_{2}^{\epsilon_{2}}\cdots a_{v}^{\epsilon_{v}},~{}~{}\epsilon_{l}=\pm 1,~{}~{}a_{l}\in\{\sigma_{1},\sigma_{2},\ldots,\sigma_{n-1},\rho^{(0)}_{1},\rho^{(0)}_{2},\ldots,\rho^{(0)}_{n-1},\ldots,\rho^{(k-1)}_{1},\rho^{(k-1)}_{2},\ldots,\rho^{(k-1)}_{n-1}\},$
the word
$\tau(u)=s_{k_{1},a_{1}}^{\epsilon_{1}}s_{k_{2},a_{2}}^{\epsilon_{2}}\cdots s_{k_{v},a_{v}}^{\epsilon_{v}},$
in the generators of $M_{k}VP_{n}$, where $k_{j}$ is the $(j-1)^{th}$ initial segment of the word $u$ if $\epsilon_{j}=1$, and a representative of the $j$-th initial segment of $u$ if $\epsilon_{j}=-1$.
By Theorem 2.9 in [18], the group $M_{k}VP_{n}$ is defined by the relations
$r_{\mu,\lambda}=\tau(\lambda r_{\mu}\lambda^{-1})=\lambda\tau(r_{\mu})\lambda^{-1},~{}~{}~{}\lambda\in\Lambda_{n},$
where $r_{\mu}$ is a defining relation of $M_{k}VB_{n}$.
Let us consider relations of $M_{k}VB_{n}$ stated in Section 3. Denote by $r_{1}=\left(\rho^{(\alpha)}_{i}\right)^{2}=1$. Then
$r_{1,e}=\tau(r_{1})$ $=s_{e,\rho^{(\alpha)}_{i}}s_{\rho_{i},\rho^{(\alpha)}_{i}}$
$=(\rho^{(\alpha)}_{i}\rho^{(0)}_{i})(\rho^{(0)}_{i}\rho^{(\alpha)}_{i}\rho^{(0)}_{i}\rho^{(0)}_{i})$
$=e.$
Denote by $r_{2}=\sigma_{i}\sigma_{j}\sigma_{i}^{-1}\sigma_{j}^{-1}$, $|i-j|\geq 2$. Then
$r_{2,e}=\tau(r_{2})$ $=s_{e,\sigma_{i}}s_{\rho_{i},\sigma_{j}}s^{-1}_{\rho_{i}\rho_{j}\rho_{i},\sigma_{i}}s^{-1}_{\rho_{i}\rho_{j}\rho_{i}\rho_{j},\sigma_{j}}$
$=(\lambda^{(0)}_{i,i+1})^{-1}(\lambda^{(0)}_{j,j+1})^{-1}(\lambda^{(0)}_{i,i+1})(\lambda^{(0)}_{j,j+1}).$
The remaining $r_{2,\lambda}$, $\lambda\in\Lambda_{n}$, can be obtained from this relation using conjugation by $\lambda^{-1}$ and it gives the same relation, by Lemma 2.1. We have obtained $\lambda_{ij}^{(0)}\lambda_{kl}^{(0)}=\lambda_{kl}^{(0)}\lambda_{ij}^{(0)}$, where indices are different.
Now, consider the next relation $r_{3}=\rho_{i}^{(\alpha)}\rho_{j}^{(\alpha)}\rho_{i}^{(\alpha)}\rho_{j}^{(\alpha)}$, $|i-j|\geq 2$. We have,
$r_{3,e}=\tau(r_{3})$ $=s_{e,\rho_{i}^{(\alpha)}}s_{\rho_{i},\rho_{j}^{(\alpha)}}s_{\rho_{i}\rho_{j},\rho_{i}^{(\alpha)}}s_{\rho_{i}\rho_{j}\rho_{i},\rho_{j}^{(\alpha)}}$
$=(\lambda_{i,i+1}^{(\alpha)})^{-1}(\lambda_{j,j+1}^{(\alpha)})^{-1}(\lambda_{i,i+1}^{(\alpha)})^{-1}(\lambda_{j,j+1}^{(\alpha)})^{-1}.$
The remaining $r_{3,\lambda}$, $\lambda\in\Lambda_{n}$. We have obtained $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\alpha)}=\lambda_{kl}^{(\alpha)}\lambda_{ij}^{(\alpha)}$, where indices are different. The other commutative relations give rises to the relations: $\lambda_{ij}^{(\alpha)}\lambda_{kl}^{(\beta)}=\lambda_{kl}^{(\beta)}\lambda_{ij}^{(\alpha)}$, $10\leq\alpha\leq\beta\leq k-$ 1. Combining all these relations, we obtain the relation (7).
Consider the next relation $r_{4}=\sigma_{i}\sigma_{i+1}\sigma_{i}\sigma^{-1}_{i+1}\sigma^{-1}_{i}\sigma^{-1}_{i+1}$.
$r_{4,e}=\tau(r_{4})$ $=s_{e,\sigma_{i}}s_{\rho_{i},\sigma_{i+1}}s_{\rho_{i}\rho_{i+1},\sigma_{i}}s^{-1}_{\rho_{i+1}\rho_{i},\sigma_{i+1}}s_{\rho_{i+1},\sigma_{i}}s^{-1}_{e,\sigma_{i+1}}$
$=(\lambda^{(0)}_{i,i+1})^{-1}(\lambda^{(0)}_{i,i+2})^{-1}(\lambda^{(0)}_{i+1,i+2})^{-1}\lambda^{(0)}_{i,i+1}\lambda^{(0)}_{i,i+2}\lambda^{(0)}_{i+1,i+2}.$
The remaining relations $r_{4,\lambda},\lambda\in\Lambda_{n}$, can be obtained from this relation using conjugation by $\lambda^{-1}$. By Lemma 2.1, we obtain (11).
Let us consider the relation $r_{5}=\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}$.
For $\alpha=0$, we get trivial relation. Now, solve for $\alpha\neq 0$.
$r_{5,e}=\tau(r_{5})=$ $s_{e,\rho^{(\alpha)}_{i}}s_{\rho_{i},\rho^{(\alpha)}_{i+1}}s_{\rho_{i}\rho_{i+1},\rho^{(\alpha)}_{i}}s^{-1}_{\rho_{i+1}\rho_{i},\rho^{(\alpha)}_{i+1}}s_{\rho_{i+1},\rho^{(\alpha)}_{i}}s^{-1}_{e,\rho^{(\alpha)}_{i+1}}$
$=(\lambda^{(\alpha)}_{i,i+1})^{-1}(\lambda^{(\alpha)}_{i,i+2})^{-1}(\lambda^{(\alpha)}_{i+1,i+2})^{-1}\lambda^{(\alpha)}_{i,i+1}\lambda^{(\alpha)}_{i,i+2}\lambda^{(\alpha)}_{i+1,i+2}.$
Conjugating this relation by all representatives from $\Lambda_{n}$, by Lemma 3.4, we obtain (12). Rest of the mixed relations of $M_{k}VB_{n}$ give rise of trivial relations. Therefore, the group $M_{k}VP_{n}$ is defined by the relations (7)-(12). ∎
We have decomposition $M_{k}VB_{n}=M_{k}VP_{n}\rtimes S_{n}$.
Corollary 3.6.
  1. (1)
    $M_{k}VB_{n}=M_{k}VP_{n}\rtimes S_{n}$.
  2. (2)
    $\langle\lambda_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\cong VP_{n}$.
  3. (3)
    For any $\alpha\in\{1,2,\ldots,k-1\}$ holds $\langle\lambda_{ij}^{(\alpha)}~{}|~{}1\leq i\not=j\leq n\rangle\cong FVP_{n}$.
  4. (4)
    If $n=3$, then $M_{k}VP_{3}\cong VP_{3}*\underbrace{(FVP_{3})*\cdots*(FVP_{3})}_{(k-1)}$.
Proof.
  • (1)
    This holds true based on the definition of $M_{k}VP_{n}$.
  • (2)
    Define a endomorphism from $\phi_{1}\colon\langle\lambda_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\to VP_{n}$ such that
    $\lambda_{ij}^{(0)}\to\lambda_{ij}.$
    Define a section of $\phi_{1}$, an inclusion map $\iota\colon VP_{n}\hookrightarrow\langle\lambda_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle$ such that $\phi_{1}\circ\iota$ is the identity map on $VP_{n}$. Since the homomorphism $\phi_{1}$ is bijective, $\langle\lambda_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\cong VP_{n}$.
  • (3)
    Let $\alpha\in\{1,2,\ldots,k-1\}$. Using the endomorphism $\phi_{2}\colon\langle\lambda_{ij}^{(\alpha)}~{}|~{}1\leq i\not=j\leq n\rangle\to FVP_{n}$ such that $\lambda_{ij}^{(\alpha)}\to\lambda_{ij}$, as in the previous case, we get the need assertion.
  • (4)
    It is easy to observe that
    $M_{k}VP_{3}=\langle\lambda_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq 3\rangle*\underbrace{\langle\lambda_{ij}^{(\alpha)}~{}|~{}1\leq i\not=j\leq 3\rangle*\cdots*\langle\lambda_{ij}^{(\alpha)}~{}|~{}1\leq i\not=j\leq 3\rangle}_{(k-1)}.$
    Using (2) and (3), we obtain the result $M_{k}VP_{3}\cong VP_{3}*\underbrace{(FVP_{3})*\cdots*(FVP_{3})}_{(k-1)}$.
3.2. Multi virtual semi–pure braid group
Define a map
$\psi_{n,k}\colon M_{k}VB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto e,~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\psi_{n,k}$ is said to be a multi semi-pure virtual braid group $M_{k}VH_{n}$.
Theorem 3.7.
$M_{k}VH_{n}$ admits a presentation with the generators
$x^{(0)}_{ij},~{}~{}1\leq i\neq j\leq n,$
$x^{(\beta)}_{ij},~{}~{}1\leq i<j\leq n,~{}~{}1\leq\beta\leq k-1,$
and the following relations:
(10) $x_{ij}^{(\alpha)}x_{kl}^{(\gamma)}=x_{kl}^{(\gamma)}x_{ij}^{(\alpha)},~{}~{}0\leq\alpha\leq\gamma\leq k-1,$
(11) $x_{ik}^{(0)}x_{kj}^{(0)}x_{ik}^{(0)}=x_{kj}^{(0)}x_{ik}^{(0)}x_{kj}^{(0)},~{}~{}1\leq i,j,k\leq n,$
(12) $x_{ik}^{(\beta)}x_{jk}^{(\beta)}x_{ij}^{(\beta)}=x_{ij}^{(\beta)}x_{jk}^{(\beta)}x_{ik}^{(\beta)},~{}~{}1\leq i<j<k\leq n,~{}~{}1\leq\beta\leq k-1,$
where distinct letters stand for distinct indices. The generators of $M_{k}VH_{n}$ can be expressed in terms of the generators of $M_{k}VB_{n}$ by the formulas
$x_{i,i+1}^{(0)}$ $=\sigma_{i}^{-1},$
$x_{i+1,i}^{(0)}$ $=\rho_{i}\,x_{i,i+1}^{(0)}\,\rho_{i}=\rho_{i}\sigma_{i}^{-1}\,\rho_{i},$
$x_{i,i+1}^{(\beta)}$ $=\rho_{i}\,\rho_{i}^{(\beta)},~{}~{}1\leq\beta\leq k-1,$
for $i=1,2,\ldots,n-1$, and
$x_{i,j}^{(\alpha)}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,x_{i,i+1}^{(\alpha)}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
$x_{j,i}^{(0)}$ $=\rho_{j-1}\,\rho_{j-2}\ldots\rho_{i+1}\,x_{i+1,i}^{(0)}\,\rho_{i+1}\ldots\rho_{j-2}\,\rho_{j-1},$
for $11\leq i<j-1\leq n-1,~{}~{}0\leq\alpha\leq k-$ 1.
Proof.
It is similar to the proof of Theorem 3.5. ∎
Corollary 3.8.
  1. (1)
    $M_{k}VB_{n}=M_{k}VH_{n}\rtimes S_{n}$.
  2. (2)
    $\langle x_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\cong VH_{n}$.
  3. (3)
    For any $\beta\in\{1,2,\ldots,k-1\}$ holds $\langle x_{ij}^{(\beta)}~{}|~{}1\leq i\not=j\leq n\rangle\cong FVP_{n}$.
  4. (4)
    If $n=3$, then $M_{k}VH_{3}\cong VH_{3}*\underbrace{(FVP_{3})*\cdots*(FVP_{3})}_{(k-1)}$.
Proof.
  • (1)
    This holds true based on the definition of $M_{k}VH_{n}$.
  • (2)
    Define a endomorphism from $\phi^{\prime}_{1}\colon\langle x_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\to VH_{n}$ such that
    $x_{ij}^{(0)}\to x_{ij}.$
    Define a section of $\phi^{\prime}_{1}$, an inclusion map $\iota\colon VH_{n}\hookrightarrow\langle x_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle$ such that $\phi^{\prime}_{1}\circ\iota$ is the identity map on $VH_{n}$. Since the homomorphism $\phi^{\prime}_{1}$ is bijective. Hence, $\langle x_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq n\rangle\cong VH_{n}$.
  • (3)
    Similar to the above proof. Let $\beta\in\{1,2,\ldots,k-1\}$. Using the endomorphism $\phi^{\prime}_{2}\colon\langle x_{ij}^{(\beta)}~{}|~{}1\leq i\not=j\leq n\rangle\to FVP_{n}$ such that $x_{ij}^{(\beta)}\to x_{ij}$.
  • (4)
    It is easy to observe that
    $M_{k}VH_{3}=\langle x_{ij}^{(0)}~{}|~{}1\leq i\not=j\leq 3\rangle*\underbrace{\langle x_{ij}^{(\beta)}~{}|~{}1\leq i\not=j\leq 3\rangle*\cdots*\langle x_{ij}^{(\beta)}~{}|~{}1\leq i\not=j\leq 3\rangle}_{(k-1)}.$
    Using (2) and (3), we obtain the result $M_{k}VH_{3}\cong VH_{3}*\underbrace{(FVP_{3})*\cdots*(FVP_{3})}_{(k-1)}$.
We have decomposition $M_{k}VB_{n}=M_{k}VH_{n}\rtimes S_{n}$ and $S_{n}$ acts on $M_{k}VH_{n}$ by the rules
Lemma 3.9.
Let $a$ be an element of $\langle\rho_{1},\rho_{2},\ldots,\rho_{n-1}\rangle$ and $\bar{a}$ is its image in $S_{n}$ under the isomorphism $\rho_{i}\mapsto(i,i+1)$, $i=1,2,\ldots,n-1$, then for any generator $x_{ij}^{(\alpha)}$ of $M_{k}VH_{n}$ the following holds
$a^{-1}x_{ij}^{(\alpha)}a=x_{(i)\bar{a},(j)\bar{a}}^{(\alpha)},$
where $(k)\bar{a}$ is the image of $k$ under the action of the permutation $\bar{a}$ and we use the relations
$x_{ji}^{(\beta)}=\left(x_{ij}^{(\beta)}\right)^{-1},~{}~{}~{}~{}1\leq\beta\leq k-1.$
3.3. Symmetric multi-virtual braid group
In the definition of $M_{k}VB_{n}$, virtual crossings need not detour over each other. The detour moves consist of only one type of virtual crossing. Also, L. Kauffman [14] suggested to study analogous of this group in which are allowed detour moves for any types of virtual crossings.
Let $k\geq 2$. Symmetric $k$-multiple virtual braid group $\widetilde{M_{k}VB}_{n}$ is a group which is quotient of $M_{k}VB_{n}$ by the forbidden relations $F3$. But in $\widetilde{M_{k}VB}_{n}$ the following relations are also forbidden:
$F1\colon\sigma_{i}\sigma_{i+1}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\sigma_{i}\sigma_{i+1}$, $10\leq\alpha\leq k-$ 1;
$F2\colon\sigma_{i+1}\sigma_{i}\,\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\,\sigma_{i+1}\sigma_{i}$, $10\leq\alpha\leq k-$ 1;
Hence, $\widetilde{M_{k}VB}_{n}$ is a group which is generated by elements
$\sigma_{i},\rho^{(\alpha)}_{i},~{}~{}i=1,2,\ldots,n-1,~{}\alpha=0,1,\ldots,k-1,$
and is defined by the following types of relations:
I. Involutivity of generators
$\left(\rho^{(\alpha)}_{i}\right)^{2}=1,~{}~{}i=1,2,\ldots,n-1,~{}\alpha=0,1,\ldots,k-1.$
II. For commutativity
– homogeneous
$\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i},\,\,\rho_{i}^{(\alpha)}\rho_{j}^{(\alpha)}=\rho_{j}^{(\alpha)}\rho_{i}^{(\alpha)},~{}~{}|i-j|\geq 2;$
– mixed ($\alpha\not=\beta$)
$\sigma_{i}\rho_{j}^{(\alpha)}=\rho_{j}^{(\alpha)}\sigma_{i},\,\,\rho_{i}^{(\alpha)}\rho_{j}^{(\beta)}=\rho_{j}^{(\beta)}\rho_{i}^{(\alpha)},~{}~{}|i-j|\geq 2.$
III. Braid relations
– homogeneous
$\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1},\,\,\,\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)},~{}~{}i=1,2,\ldots,{n-2};$
– mixed ($\alpha<\beta$)
$\sigma_{i}\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\sigma_{i+1},$
$\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)},~{}~{}~{}~{}\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\beta)}\,\rho_{i+1}^{(\alpha)},~{}~{}i=1,2,\ldots,{n-2}.$
Example 3.10.
For $M_{2}VB_{n}$, follow by Kauffman, we will use the following set of generators
$\left\{\sigma_{i},\rho_{i},\ \tau_{i}|1\leq i\leq n-1\right\}:$
and the defining relations:
$(1)\,\,\ \sigma_{i}\sigma_{i}^{-1}=\sigma_{i}^{-1}\sigma_{i}=1_{n};$
$(2)\,\,\ \sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i},\,\,|i-j|\geq 2;$
$(3)\,\,\ \sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(4)\,\,\rho_{i}^{2}=1_{n},\,\,i=1,2,\ldots,{n-1};$
$(5)\,\,\tau_{i}^{2}=1_{n},\,\,i=1,2,\ldots,{n-1};$
(6)    $\rho_{i}\tau_{j}=\tau_{j}\rho_{i},\,\,|i-j|\geq 2;$
$(7)\,\,\ \rho_{i}\rho_{j}=\rho_{j}\rho_{i},\,\,|i-j|\geq 2;$
$(8)\,\,\ \tau_{i}\tau_{j}=\tau_{j}\tau_{i},\,\,|i-j|\geq 2;$
$(9)\,\,\ \rho_{i}\rho_{i+1}\rho_{i}=\rho_{i+1}\rho_{i}\rho_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(10)\,\,\ \tau_{i}\tau_{i+1}\tau_{i}=\tau_{i+1}\tau_{i}\tau_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(11)\,\,\rho_{i}\tau_{i+1}\rho_{i}=\rho_{i+1}\tau_{i}\rho_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(12)\,\,\rho_{i}\sigma_{i+1}\rho_{i}=\rho_{i+1}\sigma_{i}\rho_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(13)\,\,\rho_{i}\rho_{i+1}\sigma_{i}=\sigma_{i+1}\rho_{i}\rho_{i+1},\,\,i=1,2,\ldots,{n-1};$
$(14)\,\,\rho_{i}\ \rho_{i+1}\tau_{i}=\tau_{i+1}\rho_{i}\rho_{i+1},\,\,i=1,2,\ldots,{n-1}.$
3.4. Symmetric multi-virtual pure braid group
Let us define a map
$\phi_{n,k}\colon~{}~{}\widetilde{M_{k}VB_{n}}\to S_{n},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\phi_{n,k}$ is the symmetric multi-virtual pure braid group $\widetilde{M_{k}VP_{n}}$.
Using the fact that $\widetilde{M_{k}VB}_{n}$ is the quotient of $M_{k}VB_{n}$ by the relations
$\sigma_{i}\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\sigma_{i+1},~{}~{}~{}~{}\text{ and }~{}~{}~{}~{}\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)},$
let us find a presentation for $\widetilde{M_{k}VP}_{n}$.
From the relation $\sigma_{i}\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\sigma_{i+1}\text{, for }\alpha\neq 0,$ follows the relation:
$\lambda_{ij}^{(0)}\lambda_{ik}^{(\alpha)}\lambda_{jk}^{(\alpha)}=\lambda_{jk}^{(\alpha)}\lambda_{ik}^{(\alpha)}\lambda_{ij}^{(0)}.$
From the relation $\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}$, for $0<\alpha<\beta$, we obtain relation:
$\lambda_{ij}^{(\alpha)}\lambda_{ik}^{(\alpha)}\lambda_{jk}^{(\beta)}=\lambda_{jk}^{(\beta)}\lambda_{ik}^{(\alpha)}\lambda_{ij}^{(\alpha)}.$
We have
Proposition 3.11.
$\widetilde{M_{k}VP}_{n}$ is the quotient of $M_{k}VP_{n}$ by the relations
$\lambda_{ij}^{(0)}\lambda_{ik}^{(\alpha)}\lambda_{jk}^{(\alpha)}=\lambda_{jk}^{(\alpha)}\lambda_{ik}^{(\alpha)}\lambda_{ij}^{(0)},~{}~{}\alpha\neq 0,$
$\lambda_{ij}^{(\alpha)}\lambda_{ik}^{(\alpha)}\lambda_{jk}^{(\beta)}=\lambda_{jk}^{(\beta)}\lambda_{ik}^{(\alpha)}\lambda_{ij}^{(\alpha)},~{}~{}0<\alpha<\beta.$
3.5. Symmetric multi-virtual semi-pure braid group
Let us define a map
$\psi_{n,k}\colon~{}~{}\widetilde{M_{k}VB_{n}}\to S_{n},~{}~{}\sigma_{i}\mapsto e,~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\psi_{n,k}$ is the symmetric multi semi-pure virtual braid group $\widetilde{M_{k}VH_{n}}$.
Let us find a presentation for $\widetilde{M_{k}VH}_{n}$. Since $\widetilde{M_{k}VB}_{n}$ it is the quotient $M_{k}VB_{n}$ by the relations
$\sigma_{i}\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\sigma_{i+1},~{}~{}~{}~{}\text{ and }~{}~{}~{}~{}\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)},$
from the relation $\sigma_{i}\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\alpha)}\sigma_{i+1}\text{, for }\alpha\neq 0,$ we obtain the relation:
$x_{ij}^{(0)}x_{ik}^{(\alpha)}x_{jk}^{(\alpha)}=x_{ik}^{(\alpha)}x_{jk}^{(\alpha)}x_{ij}^{(0)};$
from the relation $\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}$, for $0<\alpha<\beta$, we obtain relation:
$x_{ij}^{(\alpha)}x_{ik}^{(\alpha)}x_{jk}^{(\beta)}=x_{jk}^{(\beta)}x_{ik}^{(\alpha)}x_{ij}^{(\alpha)}.$
Proposition 3.12.
$\widetilde{M_{k}VH}_{n}$ is the quotient of $M_{k}VH_{n}$ by the relations
$x_{ij}^{(0)}x_{ik}^{(\alpha)}x_{jk}^{(\alpha)}=x_{ik}^{(\alpha)}x_{jk}^{(\alpha)}x_{ij}^{(0)},~{}~{}\alpha\neq 0,$
$x_{ij}^{(\alpha)}x_{ik}^{(\alpha)}x_{jk}^{(\beta)}=x_{jk}^{(\beta)}x_{ik}^{(\alpha)}x_{ij}^{(\alpha)},~{}~{}0<\alpha<\beta.$
4. Some quotients and subgroups of multi-virtual braid group
If we add to $M_{k}VB_{n}$ the forbidden relation $F1$ for all $i=1,2,\ldots,n-2,~{}\alpha=0,1,\ldots,k-1,$ we get the multi-welded braid group $M_{k}WB_{n}$. If we add to $M_{k}VB_{n}$ the forbidden relation $F1$ and $F2$ for all $i=1,2,\ldots,n-2,~{}\alpha=0,1,\ldots,k-1,$ we get the multi-unrestricted braid group $M_{k}UB_{n}$.
In the previous section we defined two maps:
$\varphi_{n,k}\colon M_{k}VB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1,$
with the kernel $M_{k}VP_{n}=\ker\varphi_{n,k}$, the multi-virtual pure braid group, and the map
$\psi_{n,k}\colon M_{k}VB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto e,~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1,$
with the kernel $M_{k}VH_{n}=\ker\psi_{n,k}$, the multi-virtual semi-pure braid group.
The first map induces the map
$\varphi_{n,k}\colon M_{k}WB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\varphi_{n,k}$ is the multi-welded pure braid group $M_{k}WP_{n}$. The second map induces the map
$\psi_{n,k}\colon M_{k}WB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto e,~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\psi_{n,k}$ is the multi-welded semi-pure braid group $M_{k}WH_{n}$.
Also, the first map induces the map
$\varphi_{n,k}\colon M_{k}UB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\varphi_{n,k}$ is the multi-unrestricted pure braid group $M_{k}UP_{n}$. The second map induces the map
$\psi_{n,k}\colon M_{k}UB_{n}\to S_{n},~{}~{}\sigma_{i}\mapsto e,~{}~{}\rho_{i}^{(\alpha)}\mapsto\rho_{i},~{}~{}i=1,2,\ldots,n-2,~{}\alpha=0,1,2,\ldots,k-1.$
The kernel $\ker\psi_{n,k}$ is the multi-unrestricted semi-pure braid group $M_{k}UH_{n}$.
The generators and defining relations for the pure braid groups $M_{k}WP_{n}$ and $M_{k}UP_{n}$, and for semi-pure $M_{k}WH_{n}$ and $M_{k}UH_{n}$ are stated as follows. The group $M_{k}WB_{n}$ is quotient of the group $M_{k}VB_{n}$ by relation $F1$. From the relation $\sigma_{i}\sigma_{i+1}\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\sigma_{i}\sigma_{i+1}$, we obtain the following relation:
$\lambda_{ij}^{(0)}\lambda_{ik}^{(0)}\lambda_{jk}^{(\alpha)}=\lambda_{jk}^{(\alpha)}\lambda_{ik}^{(0)}\lambda_{ij}^{(0)}\text{, for }\alpha\neq 0.$
$\lambda_{ij}^{(0)}\lambda_{ik}^{(0)}=\lambda_{ik}^{(0)}\lambda_{ij}^{(0)}.$
The group $M_{k}UP_{n}$ is quotient of group $M_{k}WP_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relation:
$\lambda_{jk}^{(0)}\lambda_{ik}^{(0)}\lambda_{ij}^{(\alpha)}=\lambda_{ij}^{(\alpha)}\lambda_{ik}^{(0)}\lambda_{jk}^{(0)}\text{, for }\alpha\neq 0,$
$\lambda_{jk}^{(0)}\lambda_{ik}^{(0)}=\lambda_{ik}^{(0)}\lambda_{jk}^{(0)}.$
Similarly, we can find presentation for the semi-pure braid groups $M_{k}WH_{n}$ and $M_{k}UH_{n}$.
The group $M_{k}WB_{n}$ is quotient of the group $M_{k}VB_{n}$ by relation $F1$.
From the relation $\sigma_{i}\sigma_{i+1}\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\sigma_{i}\sigma_{i+1}$, we obtain the following relations:
$x_{ik}^{(0)}x_{ji}^{(0)}x_{ki}^{(\alpha)}=x_{ij}^{(\alpha)}x_{jk}^{(0)}x_{ij}^{(0)}\text{, for }\alpha\neq 0,$
$x_{jk}^{(0)}x_{ik}^{(0)}=x_{ik}^{(0)}x_{jk}^{(0)}.$
The group $M_{k}UH_{n}$ is quotient of group $M_{k}WH_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relations:
$x_{ik}^{(0)}x_{kj}^{(0)}x_{ki}^{(\alpha)}=x_{jk}^{(\alpha)}x_{ij}^{(0)}x_{jk}^{(0)}\text{, for }\alpha\neq 0,$
$x_{jk}^{(0)}x_{ik}^{(0)}=x_{ik}^{(0)}x_{jk}^{(0)}.$
5. Three strands multi-virtual braid group
To illustrate definitions and results of the previous sections, we consider 2-virtual 3-strand braid group, some its quotients and subgroups. As rule, 3-strand braid groups are different from the braid groups with bigger number of strands and are studied separately (see [6], [8], [12], [16]).
The group $MVB_{3}=M_{2}VB_{3}$ is generated by 6 elements:
$\sigma_{1},~{}~{}\sigma_{2},~{}~{}\rho_{1},~{}~{}\rho_{2},~{}~{}\tau_{1},~{}~{}\tau_{2},$
and is defined by relations
$\rho_{1}^{2}=\rho_{2}^{2}=\tau_{1}^{2}=\tau_{2}^{2}=e,$
$\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2},~{}~{}\rho_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\rho_{2},~{}~{}\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2},$
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2},~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2}.$
It is not difficult to see that $\langle\sigma_{1},~{}~{}\sigma_{2},~{}~{}\rho_{1},~{}~{}\rho_{2}\rangle\cong VB_{3}$ is the virtual braid group and $\langle\sigma_{1},~{}~{}\sigma_{2},~{}~{}\tau_{1},~{}~{}\tau_{2}\rangle\cong FVB_{3}$ is the flat virtual braid group.
Consider a homomorphism which is defined on the generators:
$\varphi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\varphi_{3}$ is the 2-virtual pure braid group $MVP_{3}$. Let us find its presentation. Using the Reidemeister-Schreier method, we see that this group is generated by elements
$\lambda_{12}=\rho_{1}\sigma_{1}^{-1},~{}~{}\lambda_{21}=\sigma_{1}^{-1}\rho_{1},~{}~{}\lambda_{23}=\rho_{2}\sigma_{2}^{-1},~{}~{}\lambda_{32}=\sigma_{2}^{-1}\rho_{2},$
$\lambda_{13}=\rho_{2}(\rho_{1}\sigma_{1}^{-1})\rho_{2}=\rho_{2}\lambda_{12}\rho_{2},~{}~{}\lambda_{31}=\rho_{2}(\sigma_{1}^{-1}\rho_{1})\rho_{2}=\rho_{2}\lambda_{21}\rho_{2}.$
$\mu_{12}=\rho_{1}\tau_{1},~{}~{}\mu_{21}=\tau_{1}\rho_{1},~{}~{}\mu_{23}=\rho_{2}\tau_{2},~{}~{}\mu_{32}=\tau_{2}\rho_{2},$
$\mu_{13}=\rho_{2}(\rho_{1}\tau_{1})\rho_{2}=\rho_{2}\mu_{12}\rho_{2},~{}~{}\mu_{31}=\rho_{2}(\tau_{1}\rho_{1})\rho_{2}=\rho_{2}\mu_{21}\rho_{2}.$
From the relations
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2},~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2}.$
follows
$\lambda_{12}^{\rho_{2}\rho_{1}}=\lambda_{23},~{}~{}~{}\mu_{12}^{\rho_{2}\rho_{1}}=\mu_{23}$
and all other formulas of conjugations by elements $\rho_{i}$.
From relations $\tau_{1}^{2}=\tau_{2}^{2}=1$ follow relations
$\mu_{21}=\mu_{12}^{-1},~{}~{}\mu_{32}=\mu_{23}^{-1},~{}~{}\mu_{31}=\mu_{13}^{-1}.$
Hence, we can remove elements $\mu_{21}$, $\mu_{32}$ and $\mu_{31}$ from the list of generators $MVP_{3}$.
From relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ using the rewritable process and formulas of conjugations by elements $\rho_{i}$ we get 6 relations:
$\lambda_{12}\lambda_{32}\lambda_{13}=\lambda_{13}\lambda_{32}\lambda_{12},~{}~{}~{}\lambda_{21}\lambda_{31}\lambda_{23}=\lambda_{23}\lambda_{31}\lambda_{21},$
$\lambda_{13}\lambda_{23}\lambda_{12}=\lambda_{12}\lambda_{23}\lambda_{13},~{}~{}~{}\lambda_{31}\lambda_{21}\lambda_{32}=\lambda_{32}\lambda_{21}\lambda_{31},$
$\lambda_{23}\lambda_{13}\lambda_{21}=\lambda_{21}\lambda_{13}\lambda_{23},~{}~{}~{}\lambda_{32}\lambda_{12}\lambda_{31}=\lambda_{31}\lambda_{12}\lambda_{32}.$
From relation $\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2}$ we do not have new relations.
From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows relation
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
Hence, we have
Proposition 5.1.
The group $MVP_{3}$ is generated by elements
$\lambda_{12},~{}~{}\lambda_{21},~{}~{}\lambda_{13},~{}~{}\lambda_{31},~{}~{}\lambda_{23},~{}~{}\lambda_{32},~{}~{}\mu_{12},~{}~{}\mu_{13},~{}~{}\mu_{23}$
and is defined by the relations
$\lambda_{12}\lambda_{32}\lambda_{13}=\lambda_{13}\lambda_{32}\lambda_{12},~{}~{}~{}\lambda_{21}\lambda_{31}\lambda_{23}=\lambda_{23}\lambda_{31}\lambda_{21},$
$\lambda_{13}\lambda_{23}\lambda_{12}=\lambda_{12}\lambda_{23}\lambda_{13},~{}~{}~{}\lambda_{31}\lambda_{21}\lambda_{32}=\lambda_{32}\lambda_{21}\lambda_{31},$
$\lambda_{23}\lambda_{13}\lambda_{21}=\lambda_{21}\lambda_{13}\lambda_{23},~{}~{}~{}\lambda_{32}\lambda_{12}\lambda_{31}=\lambda_{31}\lambda_{12}\lambda_{32},$
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
5.1. 2-virtual semi-pure braid group on 3 strands
In this subsection we consider the map
$\psi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\psi_{3}$ is the 2-virtual semi-pure braid group $MVH_{3}$. Let us find its representation. This group is generated by elements
$x_{12}=\sigma_{1},~{}~{}x_{23}=\sigma_{2},~{}~{}x_{13}=\rho_{2}\sigma_{1}\rho_{2},$
$x_{21}=\rho_{1}\sigma_{1}\rho_{1},~{}~{}x_{32}=\rho_{2}\sigma_{2}\rho_{2},~{}~{}x_{31}=\rho_{2}\rho_{1}\sigma_{1}\rho_{1}\rho_{2},$
$z_{12}=\rho_{1}\tau_{1},~{}~{}z_{23}=\rho_{2}\tau_{2},~{}~{}z_{13}=\rho_{2}\rho_{1}\tau_{1}\rho_{2},$
$z_{21}=\tau_{1}\rho_{1},~{}~{}z_{32}=\tau_{2}\rho_{2},~{}~{}z_{31}=\rho_{2}\tau_{1}\rho_{1}\rho_{2}.$
Further, from relation $\tau_{i}^{2}=e$ follows relations $z_{21}z_{12}=z_{31}z_{13}=z_{32}z_{23}=e$.
From the relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ follows 6 relation
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}.$
From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows relation
$z_{12}z_{13}z_{23}=z_{23}z_{13}z_{12}.$
The relations
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2},~{}~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2}$
give only conjugations rules by elements $\rho_{i}$.
Proposition 5.2.
  1. (1)
    The subgroup of $MVH_{3}$ which is generated by elements $x_{ij}$ is isomorphic to $VH_{3}$.
  2. (2)
    $\langle z_{12},z_{13},z_{23}~{}|~{}z_{12}z_{13}z_{23}=z_{23}z_{13}z_{12}\rangle\cong FVP_{3}$.
  3. (3)
    $MVH_{3}\cong VH_{3}*FVP_{3}$.
5.2. Some other maps
In this subsection we consider the map
$\chi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
Denote its kernel $\ker\chi_{3}$ by $MVQ_{3}$ and find its presentation. It is easy to see that this group is generated by elements
$x_{12}=\sigma_{1},~{}~{}x_{23}=\sigma_{2},~{}~{}x_{13}=\rho_{2}\sigma_{1}\rho_{2},$
$x_{21}=\rho_{1}\sigma_{1}\rho_{1},~{}~{}x_{32}=\rho_{2}\sigma_{2}\rho_{2},~{}~{}x_{31}=\rho_{2}\rho_{1}\sigma_{1}\rho_{1}\rho_{2},$
$y_{12}=\tau_{1},~{}~{}y_{23}=\tau_{2},~{}~{}y_{13}=\rho_{2}\tau_{1}\rho_{2},$
$y_{21}=\rho_{1}\tau_{1}\rho_{1},~{}~{}y_{32}=\rho_{2}\tau_{2}\rho_{2},~{}~{}y_{31}=\rho_{2}\rho_{1}\tau_{1}\rho_{1}\rho_{2}.$
Further, from the relation $\tau_{i}^{2}=e$ follows the relations $y_{ij}^{2}=e$; from the relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ follows 6 relations:
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13};$
from the relations $\rho_{1}\rho_{2}\sigma_{1}=\sigma_{2}\rho_{1}\rho_{2}$ and $\rho_{1}\rho_{2}\tau_{1}=\tau_{2}\rho_{1}\rho_{2}$ we do not have new relations; from the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ we have relations
$y_{12}y_{23}y_{12}=y_{23}y_{12}y_{23},~{}~{}y_{12}y_{13}y_{12}=y_{13}y_{12}y_{13},~{}~{}y_{13}y_{23}y_{13}=y_{23}y_{13}y_{23}.$
Proposition 5.3.
The group is $MVQ_{3}$ is generated by elements
$x_{12},~{}~{}x_{23},~{}~{}x_{13},~{}~{}x_{21},~{}~{}x_{32},~{}~{}x_{31},~{}~{}y_{12},~{}~{}y_{13}.~{}~{}y_{23},$
and is defined by relations
$(y_{12}y_{23})^{3}=(y_{12}y_{13})^{3}=(y_{23}y_{13})^{3}=e,$
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}.$
Corollary 5.4.
$\langle x_{12},~{}~{}x_{23},~{}~{}x_{13},~{}~{}x_{21},~{}~{}x_{32},~{}~{}x_{31}\rangle\cong H_{1}*H_{2},$
where
$H_{1}=\langle x_{13},~{}~{}x_{21},~{}~{}x_{32}~{}|~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},~{}~{}x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}\rangle,$
$H_{2}=\langle x_{12},~{}~{}x_{31},~{}~{}x_{23}~{}|~{}x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},~{}~{}x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31}\rangle.$
Corollary 5.5.
Let
$Y=\langle y_{12},~{}~{}y_{13},~{}~{}y_{23}~{}|~{}y_{12}^{2}=y_{13}^{2}=y_{23}^{2}=(y_{12}y_{13})^{3}=(y_{12}y_{23})^{3}=(y_{23}y_{13})^{3}=e\rangle.$
Hence, $MVQ_{3}\cong H_{1}*H_{2}*Y.$
Consider the map
$\chi_{4}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\chi_{4}$ is denoted as $MVC_{3}$. It is easy to check
Proposition 5.6.
$MVC_{3}=VP_{3}*Y$.
5.3. Symmetric 2-virtual braid group on 3 strands
This group $\widetilde{MVB}_{3}$ is generated by 6 elements:
$\sigma_{1},~{}~{}\sigma_{2},~{}~{}\rho_{1},~{}~{}\rho_{2},~{}~{}\tau_{1},~{}~{}\tau_{2},$
and is defined by relations
$\rho_{1}^{2}=\rho_{2}^{2}=\tau_{1}^{2}=\tau_{2}^{2}=1,$
$\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2},~{}~{}\rho_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\rho_{2},~{}~{}\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2},$
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{1}\rho_{2}\sigma_{2},~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2},$
and extra mixed relations
(13) $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2},~{}~{}\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}.$
Let us defined a map on the generators
$\varphi_{3}\colon\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\varphi_{3}$ is the symmetric 2-virtual pure braid group $\widetilde{MVP}_{3}$. Let us find its presentation. Since it is the quotient of $MVP_{3}$ by the relations (13), we have rewrite these relations in the generators of $MVP_{3}$. From the relation $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2}$ follows relation
$\lambda_{12}\,\mu_{13}\,\mu_{23}=\mu_{23}\,\mu_{13}\,\lambda_{12},~{}~{}\lambda_{21}\,\mu_{23}\,\mu_{13}=\mu_{13}\,\mu_{23}\,\lambda_{21},$
$\lambda_{23}\,\mu_{13}\,\mu_{12}=\mu_{12}\,\mu_{13}\,\lambda_{23},~{}~{}\lambda_{32}\,\mu_{12}\,\mu_{13}=\mu_{13}\,\mu_{12}\,\lambda_{32},$
$\mu_{12}\,\lambda_{13}\,\mu_{23}=\mu_{23}\,\lambda_{13}\,\mu_{12},~{}~{}\lambda_{31}\,\mu_{12}^{-1}\,\mu_{23}=\mu_{23}\,\mu_{12}^{-1}\,\lambda_{31}.$
From the relation $\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}$ follows that
$\langle\mu_{12},\mu_{13},\mu_{23}\rangle\cong\mathbb{Z}^{3}$
is a free abelian group of rank 3. From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows the relation
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
5.4. Symmetric 2-virtual semi-pure braid group
Let
$\psi_{3}\colon\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\psi_{3}$ is the symmetric 2-virtual semi-pure braid group $\widetilde{MVH}_{3}$.
The group $\widetilde{MVH}_{3}$ is the quotient $MVH_{3}$ by the relations (13), we have rewritten these relations in the generators of $MVH_{3}$. From the relation $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2}$ follows the relations
$x_{21}z_{13}z_{23}=z_{13}z_{23}x_{21},~{}~{}x_{12}z_{23}z_{13}=z_{23}z_{13}x_{12},$
$x_{31}z_{12}z_{32}=z_{12}z_{32}x_{31},~{}~{}x_{32}z_{21}z_{31}=z_{21}z_{31}x_{32},$
$x_{13}z_{32}z_{12}=z_{32}z_{12}x_{13},~{}~{}x_{23}z_{31}z_{21}=z_{31}z_{21}x_{23}.$
From the relation $\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}$ follows that
$\langle z_{12},z_{13},z_{23}\rangle\cong\mathbb{Z}^{3}$
is a free abelian group of rank 3.
Remark 5.7.
If we consider a map
$\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2,$
then from defining relations of $\widetilde{MVB}_{3}$ follows that $\rho_{1}=\rho_{2}$. Therefore this map does not form an epimorphism.
Concluding Remarks
This work opens several directions for future research within the framework of multi-virtual braid theory. Below are some open problems and potential avenues for further investigation:
  1. (1)
    Construct representations of the multi-virtual braid group using automorphisms of groups and quandles.
  2. (2)
    Investigate which algebraic structures correspond to braid-like groups. For instance, it is well known that the Hecke algebra corresponds to the classical braid group $B_{n}$. What analogous structures, if any, correspond to multi-virtual braid groups?
  3. (3)
    Define a group and a quandle for multi-virtual knots and establish that these structures are indeed invariants under appropriate equivalence moves.
  4. (4)
    Loday and Stein [17] introduced parametric braid groups. Is it possible generalized this construction to define multi-virtual braid groups?
Acknowledgments
This work has been supported by the grant the Russian Science Foundation, RSF 24-21-00102, https://rscf.ru/project/24-21-00102/.